Verify Cayley-Hamilton theorem of A= ( 1 2 2 3 1 0 1 1 1 ) 


1
Expert's answer
2021-08-20T04:25:53-0400
"A=\begin{pmatrix}
 1 & 2 & 2 \\
 3 & 1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}"

"P_A(\lambda)=\det(A-\lambda I)=\begin{vmatrix}
 1-\lambda & 2 & 2 \\
 3 & 1-\lambda & 0 \\
 1 & 1 & 1-\lambda \\
\end{vmatrix}"

"=(1-\lambda)\begin{vmatrix}
 1-\lambda & 0 \\
 1 & \lambda-1
\end{vmatrix}-2\begin{vmatrix}
 3 & 0 \\
 1 & 1-\lambda
\end{vmatrix}"

"+2\begin{vmatrix}
 3 & 1-\lambda \\
 1 & 1
\end{vmatrix}=(1-\lambda)^3-6(1-\lambda)+6-2(1-\lambda)"

"=1-3\lambda+3\lambda^2-\lambda^3-8+8\lambda+6"

"=-\lambda^3+3\lambda^2+5\lambda-1"

"A^2=\begin{pmatrix}
 1 & 2 & 2 \\
 3 & 1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}\begin{pmatrix}
 1 & 2 & 2 \\
 3 & 1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 1+6+2 & 2+2+2 & 2+0+2 \\
 3+3+0 & 6+1+0 & 6+0+0 \\
 1+3+1 & 2+1+1 & 2+0+1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 9 & 6 & 4 \\
 6 & 7 & 6 \\
 5 & 4 & 3 \\
\end{pmatrix}"

"A^3=\begin{pmatrix}
 9 & 6 & 4 \\
 6 & 7 & 6 \\
 5 & 4 & 3 \\
\end{pmatrix}\begin{pmatrix}
 1 & 2 & 2 \\
 3 & 1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 9+18+4 & 18+6+4 & 18+0+4 \\
 6+21+6 & 12+7+6 & 12+0+6 \\
 5+12+3 & 10+4+3 & 10+0+3 \\
\end{pmatrix}"

"=\begin{pmatrix}
 31 & 28 & 22 \\
 33 & 25 & 18 \\
 20 & 17 & 13 \\
\end{pmatrix}"

"-A^3+3A^2+5A-I=-\begin{pmatrix}
 31 & 28 & 22 \\
 33 & 25 & 18 \\
 20 & 17 & 13 \\
\end{pmatrix}"

"+3\begin{pmatrix}
 9 & 6 & 4 \\
 6 & 7 & 6 \\
 5 & 4 & 3 \\
\end{pmatrix}+5\begin{pmatrix}
 1 & 2 & 2 \\
 3 & 1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}-\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 -31+27+5-1 & -28+18+10-0 & -22+12+10-0 \\
 -33+18+15-0 & -25+21+5-1 & -18+18+0-0 \\
 -20+15+5-0 & -17+12+5-0 & -13+9+5-1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 0 & 0 & 0 \\
 0 & 0 & 0 \\
 0 & 0 & 0 \\
\end{pmatrix}"

The matrix "A" satisfies its own characteristic equation. 

Therefore we verify  Cayley-Hamilton theorem of "A=\begin{pmatrix}
 1 & 2 & 2 \\
 3 & 1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}."